# Multi-Partite Output Regulation of Multi-Agent Systems

### v1

@@ V Illustrative Numerical Examples
[…]
Figure 2: Signed directed graph $\mathcal{G}(\mathcal{A})$.
Except for $v^{T}(0)=[1,0]$, all the initial states are taken as 00 in the simulation. We may infer from Fig. 3 that each follower belonging to the same set in the partition tracks the $k$-partition transformation term multiple of the reference in the respected subplots, as promised by Theorem 1 and Corollary 1 if each subplot is regarded as separate MORPs. Note that the MORP reduces to the CORP in the first and third subplots.
Figure 3: The trajectories for respective time intervals. Here, $y_{ij}$ denotes the $j$th entry of follower $i$’s output or $j$th entry of $v$ while “o” marks $y_{ij}$ for each agent at $t=0$, $t=15$, and $t=35$ seconds in the respective subplots.
Example 3
This example compares the partition-dependent steps of both design strategies in terms of scalability. In particular, Fig. 4 shows the total elapsed times with an average laptop for follower 2 in Example 2 to complete steps (iii) and (iv) of the first design (see Remark 6) and step (v) of the second design (see Remark 11) as the cardinality of the given set [^6] Such sets are generated using $randn$ function in MATLAB. of $k$-partition transformation terms increases. With the first design strategy, feedforward gains for up to $270$ $k$-partition transformation terms can be computed within $6$ milliseconds. On the other hand, with the second one, feedforward gains for approximately $10000$ $k$-partition transformation terms can be computed within the same amount of time.
@@ VI Conclusion
The MORP for linear MASs has been formulated and solved for the first time. The primary motivation has been to provide MASs with objectives beyond cooperation and bipartition for tactical flexibility in adverse operating environments.

### v2

@@ V Experimental and Numerical Illustrations
This example compares the partition-dependent steps of both design strategies in terms of scalability. In particular, Fig. 3 shows the total elapsed times with an average laptop for follower 1 in Example 1 with $D_{1}=[0,1]$ and $E^{T}_{1}=[0,-4]$ to complete steps (iii) and (iv) of the first design (see Remark 6) and step (v) of the second design (see Remark 11) as the cardinality of the given set [^9] Such sets are generated using $randn$ function in MATLAB. of $k$-partition transformation terms increases. With the first design strategy, feedforward gains for up to $190$ $k$-partition transformation terms can be computed within $4$ milliseconds. On the other hand, with the second one, feedforward gains for approximately $10000$ $k$-partition transformation terms can be computed within the same amount of time.
@@ VI Conclusion
The primary motivation of this paper has been to provide MASs with objectives beyond cooperation and bipartition for tactical flexibility in adverse operating environments. To this end, the MORP for linear MASs has been formulated and solved for the first time. Two design strategies for the control parameters have been proposed. The first applies to a broader set of MASs, but it has a drawback due to the partition-dependent regulator equations. The second eliminates this drawback, and hence, it is significantly more scalable, yet applicable only when the followers are subject to matched disturbances. Table I summarizes the differences between the first and second design strategies. Theoretical results have been demonstrated by experimental and numerical tests.
The distributed observer in (III-A) assumes all followers have access to the matrix $A_{0}$. To relax this assumption to a small subset of the followers, solving the MORP with distributed control laws involving adaptive distributed observers (e.g., see [7]) would be well worth an exploration. To widen the application domain of the MORP, another research direction is investigating $k$-partition transformation generation as a distributed optimization problem, where the followers optimize a MAS-level objective such as total fuel or energy consumption.